OpenStudy (anonymous):

Find f ''(x) when f(x)=sinx/ x^2. Thanks

OpenStudy (anonymous):

i suggest graphing it to get your 0s

OpenStudy (anonymous):

I dont know how to do that though.

OpenStudy (anonymous):

:/ Pleasee help me with this. Ive been working on this problem for like an hour and I still can't get it. Ugh im so confused

OpenStudy (anonymous):

Rewrite this as \[\sin(x) \times x^{-2}\] Then use the rule of taking the 1st times the derivative of the 2nd plus the 2nd times the derivative of the first. (1D2 + 2D1 is what I like to remember it by). \[\sin(x)(-2x ^{-3)}+(x ^{-2})(-\cos(x))\] is the first derivative. Do the same for the second derivative now using the 1D2 + 2D1 rule for both terms.

OpenStudy (anonymous):

WHATTT?! please break it down more for me! :/

OpenStudy (anonymous):

got it!

OpenStudy (anonymous):

You sure? :)

OpenStudy (anonymous):

yuppp :D You helped me alot! you're the best <3